### v3

# Comments, corrections, and related references welcomed, as always! ed October 26, 2025

Abstract.
If $\mathcal{C}$ is a category of algebras closed under finite direct products, and $M_{\mathcal{C}}$ the commutative monoid of isomorphism classes of members of $\mathcal{C},$ with operation induced by direct product, A. Tarski defined a nonidentity element $p$ of $M_{\mathcal{C}}$ to be prime if, whenever it divides a product of two elements, it divides one of them, and defined an object of $\mathcal{C}$ to be prime if its isomorphism class has this property.
McKenzie, McNulty and Taylor [9, p. 263] ask whether the category of nonempty semigroups has any objects that are prime in this sense. We show in §2 that it does not. However, for the category of monoids, and some other subcategories of semigroups, we obtain examples of prime objects in §§3-4. In §5 two related questions from [9], open so far as I know, are recalled.
In §6, which can be read independently of the rest of this note, we recall two conditions called primeness by semigroup theorists, and obtain results and examples on the relationships among those conditions and Tarski’s in categories of groups. §7 notes a characterization of one of those conditions on finite algebras in an arbitrary variety.
Several questions are raised.
Key words and phrases:
semigroups prime in the sense of Tarski; groups prime in senses of Tarski and of Rhodes
2020 Mathematics Subject Classification:
Primary: 20D40, 20M10. Secondary: 08A10, 18A32, 20E34.
Archived at http://arxiv.org/abs/2409.15541 . Readable at http://math.berkeley.edu/~gbergman/papers/.
After publication, any updates, errata, related references, etc., found will be noted at http://math.berkeley.edu/~gbergman/papers/abstracts .
1. Note on terminology
As noted in the abstract, Tarski’s definition of “prime” is different from a pair of senses currently common in semigroup theory. In §6 below, the latter two conditions will be recalled, and their relationship with each other and with Tarski’s on groups examined. Until that section, prime algebra will be understood in Tarski’s sense, as indicated in Definition 2.1(ii) below.
2. The category of nonempty semigroups has no prime objects
Here are two general usages that we will follow in §§2-5:
Definition 2.1.
(i) As in [9], “semigroup” will be understood to mean “nonempty semigroup”. The category of all such semigroups will be denoted $\mathbf{Semigp}.$
(ii) Also as in [9] (p. 263, top paragraph), if $\mathcal{C}$ is a category admitting finite direct products (e.g., $\mathbf{Semigp}),$ an object $X$ of $\mathcal{C}$ other than the final object (the product of the empty family, corresponding to the identity element of the monoid of isomorphism classes; in varieties of algebras, the 1\!1\!-element algebra) will be called prime if, whenever an object $Y$ of $\mathcal{C}$ admits $X$ as a direct factor, and $Y=Y_{0}\times Y_{1}$ is another direct product decomposition of $Y,$ then one of $Y_{0},$ $Y_{1}$ admits $X$ as a direct factor.
Here are some further bits of language and notation that will be used in the present section.
Definition 2.2.
(i) By a null semigroup we shall mean a semigroup in which all pairs of elements have the same product. For any cardinal $\kappa,$ the null semigroup of cardinality $\kappa\!+\!1$ (i.e., having exactly $\kappa$ elements which are not products, in addition to the one which is) will be denoted $\textup{Null}(\kappa).$
(ii) Two elements $x$ and $x^{\prime}$ of a semigroup $S$ will be called action equivalent if for all $y\in S,$ $x\,y=x^{\prime}\,y$ and $y\,x=y\,x^{\prime}.$ (Clearly, this is an equivalence relation on $S.)$
(iii) An element $x$ of a semigroup $S$ will be called a product element if it can be written x=yzx=y\,z for $y,z\in S,$ and a non-product element otherwise.
(iv) Given a semigroup $S,$ a skeleton $S^{\prime}$ of $S$ will mean any subsemigroup of $S$ which consists of all the product elements, and exactly one representative of each action equivalence class that contains no product elements.
The definition of “skeleton” may look strange; a more natural characterization is that it is a minimal subsemigroup that contains at least one element from every action equivalence class. This guarantees that it contains every product element xy,x\,y, since given $x$ and $y,$ it will contain some $x^{\prime}$ action-equivalent to $x$ and some $y^{\prime}$ action-equivalent to $y,$ and hence will contain their product, $x^{\prime}y^{\prime}=xy^{\prime}=xy.$ But since descending chains of nonempty sets can have empty intersection, the existence of such minimal subsemigroups is not self-evident, while the existence of $S^{\prime}$ as in (iv) above is.
Clearly,
(2.1) For $S$ a semigroup, all skeleta of $S$ are isomorphic to one another.
We also observe that
(2.2) A semigroup $S$ is null if and only if it has one and only one action equivalence class.
Indeed, $\Rightarrow$ is clear. Conversely, if all elements are action equivalent, then for any $x,\,x^{\prime},y,\,y^{\prime}\in S$ we have $x^{\prime}y^{\prime}=x\,y^{\prime}=x\,y,$ showing that all products are equal.
It is also easy to see that if $S$ and $T$ are semigroups, then
(2.3) Elements $(s,t),~(s^{\prime},t^{\prime})\in S\times T$ are action equivalent if and only if $s,s^{\prime}\in S$ are action equivalent and $t,t^{\prime}\in T$ are action equivalent.
(2.4) An element $(s,t)\in S\times T$ is a product element if and only if $s\in S$ is a product element and $t\in T$ is a product element.
The next lemma gives lots of isomorphisms among products of semigroups, which will be a tool in proving the nonexistence of prime semigroups.
Lemma 2.3.
If $S$ and $T$ are semigroups having isomorphic skeleta, and $\kappa$ is an infinite cardinal greater than or equal to the cardinality of every action equivalence class in $S$ and of every action equivalence class in $T,$ then $S\times\textup{Null}(\kappa)\cong T\times\textup{Null}(\kappa).$
Moreover, every action equivalence class in that product semigroup has cardinality $\kappa,$ and contains $\kappa$ non-product elements.
Proof.
We can assume without loss of generality that $S$ and $T$ have a common skeleton; hence, in particular, that their sets of product elements are the same. So, in view of (2.4), $S\times\textup{Null}(\kappa)$ and $T\times\textup{Null}(\kappa)$ have the same product elements.
Since the common skeleton of $S$ and $T$ contains at least one representative of each action equivalence class, and since the behavior of the semigroup operation on an element is determined by which action equivalence class it belongs to, we can get an isomorphism between $S\times\textup{Null}(\kappa)$ and $T\times\textup{Null}(\kappa)$ if, for each action equivalence class of the common skeleton of $S$ and $T,$ we can define a bijection between the sets of non-product elements of the corresponding action equivalence classes of $S\times\textup{Null}(\kappa)$ and $T\times\textup{Null}(\kappa).$
But each of those non-product sets has cardinality $\kappa.$ E.g., in $S\times\textup{Null}(\kappa),$ such a set has cardinality at most $\kappa$ because it is contained in the direct product of a $\leq\kappa\!$-element action equivalence class of $S$ and the $\kappa\!+\!1\!$-element semigroup $\textup{Null}(\kappa),$ and since $\kappa$ is infinite, $\kappa(\kappa\!+\!1)=\kappa;$ while it has at least $\kappa$ elements occurring as pairs $(s,\alpha),$ where $s$ is any member of the given equivalence class of $S,$ and $\alpha$ ranges over the $\kappa$ non-product elements of $\textup{Null}(\kappa).$ Hence the desired bijections can be chosen. ∎
We can now prove a large class of cases of our desired result:
Lemma 2.4.
No non-null semigroup is prime in $\mathbf{Semigp}.$
Proof.
Let $S$ be a non-null semigroup. Take an infinite cardinal $\kappa$ larger than the cardinality of any action equivalence class of $S,$ and form a semigroup $S^{+}\supset S$ by attaching to one arbitrarily chosen action equivalence class, $A\subset S,$ $\kappa$ additional elements. (I.e., add $\kappa$ additional elements to the underlying set of $S,$ and define multiplication on the resulting set so that all products involving those elements are the same as the products one gets by putting members of $A$ in their place. In particular, the product of two of the added elements will be the same as the product of any two members of $A.)$
By Lemma 2.3, we have
(2.5) $S\times\textup{Null}(\kappa)~\cong~S^{+}\times\textup{Null}(\kappa).$
If $S$ were prime, it would be isomorphic to a direct factor of one of the factor semigroups on the right-hand side of (2.5). That factor cannot be $\textup{Null}(\kappa),$ because $S,$ being non-null, has more than one action equivalence class. On the other hand, if it were $S^{+},$ so that we could write
(2.6) $S^{+}~\cong~S\times T,$
then since $S$ has no action equivalence classes of cardinality $\kappa,$ but $S^{+}$ does, the factor $T$ must have one; but the products of that class with two different action equivalence classes in $S$ would give two action equivalence classes in $S^{+}$ both of cardinality $\kappa,$ contradicting our construction of $S^{+}.$ So $S$ is not prime. ∎
To show that semigroups $\textup{Null}(\kappa)$ are also non-prime, we need a criterion for when a semigroup can be written as a direct product with $\textup{Null}(\kappa)$ as one of the factors.
Lemma 2.5.
Let $\kappa$ be a cardinal (finite or infinite), and $S$ a semigroup. Then the following two conditions are equivalent:
(i) There exists a semigroup $S_{0}$ such that $S~\cong~S_{0}\times\textup{Null}(\kappa).$
(ii) Each action equivalence class $A$ of $S$ can be written as a disjoint union, $A=A_{0}\cup A_{1},$ such that $\textup{card}(A_{1})=\kappa~\textup{card}(A_{0}),$ and all product elements of $S$ lying in $A$ belong to $A_{0}.$
Moreover, if $S$ and $\kappa$ are finite, (ii) can be rephrased:
($i\!{}^{\prime}\!$) Each action equivalence class $A$ of $S$ has cardinality $\kappa\!+\!1$ times an integer greater than or equal to the number of product elements in $A.$
Proof.
If (i) holds, let us assume without loss of generality that $S=S_{0}\times\textup{Null}(\kappa).$ Note that for each $s\in S_{0},$ the $\kappa\!+\!1$ elements $(s,x)$ $(x\in\textup{Null}(\kappa))$ are action equivalent, and at most one (the one with $x$ the product element of $\textup{Null}(\kappa))$ can be a product element, so taking the union of these families over the elements $x$ of an action equivalence class of $S_{0},$ and writing $A_{0}$ for the set of $(s,x)$ in this union with $x$ the product element of $\textup{Null}(\kappa),$ and $A_{1}$ for the set of $(s,x)$ with non-product $x,$ we get a decomposition of the induced action equivalence class of $S$ as in (ii).
Conversely, if we have a decomposition of each action equivalence class of $S$ as in (ii), let $S_{0}$ be the union over these classes $A$ of their chosen subsets $A_{0}.$ This is easily seen to be a subsemigroup of $S$ such that $S\cong S_{0}\times\textup{Null}(\kappa).$
[…]
As promised, we can now prove
Lemma 2.6.
No semigroup $\textup{Null}(\kappa)$ is prime in $\mathbf{Semigp}.$
Proof.
[…]
If $\kappa$ is infinite, let $S$ consist of elements $x_{ij}$ $(i\in\kappa,~j\in\kappa\!+\!1),$ $y_{i}$ $(i\in\kappa)$ and $z,$ with multiplication given by
(2.7) $x_{ij}\,x_{ij^{\prime}}~=~y_{i}$ for $i\in\kappa,~j,j^{\prime}\in\kappa\!+\!1,$ while all products not of this form have value $z.$
Since all 3\!3\!-fold products equal $z,$ $S$ is associative. Note that the action equivalence classes of $S$ are of two sorts: on the one hand, the sets
(2.8) $\{x_{ij}\,|\,j\in\kappa\!+\!1\},$ one for each $i\in\kappa,$
and one more equivalence class, comprising the remaining elements:
(2.9) $\{y_{i}\,|\,i\in\kappa\}\,\cup\,\{z\}.$
Of these, the classes (2.8) clearly satisfy the condition of Lemma 2.5(ii), with $A_{0}$ the singleton $\{x_{i0}\},$ and $A_{1}$ consisting of the remaining $\kappa$ elements, but the class (2.9) does not, since all of its elements are products.
Now take any positive integer $n$ (e.g., $n=1).$ Since $\kappa$ is infinite, the finite semigroup $\textup{Null}(n)$ cannot have $\textup{Null}(\kappa)$ as a direct factor. But in the product semigroup $S\times\textup{Null}(n),$ we have effectively adjoined $n\,\kappa=\kappa=\kappa^{2}$ non-product elements $(s,i)$ $(i\in n)$ to each of the $\kappa\!-$element action equivalence classes (2.8) and (2.9) of $S;$ so the product semigroup satisfies Lemma 2.5(ii), hence has $\textup{Null}(\kappa)$ as a direct factor, giving the desired example.
Finally, suppose $\kappa$ is finite.
If $\kappa=0,$ then $\textup{Null}(\kappa)$ is the 1\!1\!-element semigroup, corresponding to the identity element of the monoid of isomorphism classes, which by definition is not prime.
If $\kappa>0,$ let $S$ be the semigroup of $2(\kappa\!+\!1)$ elements,
(2.10) $x_{i}$ $(i\in\kappa\!+\!1),$ $y_{i}$ $(i\in\kappa),$  and  $z,$
with multiplication
(2.11) $x_{i}\,x_{i^{\prime}}~=~y_{0},\,$ and all other products equal to  $z.$
Here there are two action equivalence classes, each of cardinality $\kappa\!+\!1,$ namely $\{x_{i}\}$ and $\{y_{i}\}\,\cup\,\{z\}.$ Analogously to the infinite case, we see that the first satisfies Lemma 2.5(ii), but the second does not, in this case because it has two product elements, $y_{0}$ and $z,$ so that any decomposition as in Lemma 2.5(ii) would require it to have at least $2(\kappa\!+\!1)$ elements, while it has only $\kappa\!+\!1.$ So $\textup{Null}(\kappa)$ is not a direct factor of $S.$
[…]
Combining Lemmas 2.4 and 2.6, we get
Theorem 2.7.
The category $\mathbf{Semigp}$ of (nonempty) semigroups has no prime objects.∎
Note that in proving that arbitrary semigroups are not prime, the auxiliary semigroups that we used – mainly of the form $\textup{Null}(\kappa),$ but also a few of other sorts, e.g., the $S$ in the proof of Lemma 2.6 – were all commutative, and had a zero element (an element $z$ satisfying $zx=xz=z$ for all $x).$
Hence the same reasoning shows that there are no prime objects in the category of all commutative semigroups, or in the category of all semigroups with zero element. More generally,
Corollary 2.8 (to proof).
Let $C$ be any full subcategory of $\mathbf{Semigp}$ which is closed under taking pairwise direct products and direct factors, and contains all semigroups $S$ that have a zero element $z\in S$ such that $s_{1}\,s_{2}\,s_{3}=z$ for all $s_{1},\,s_{2},\,s_{3}\in S.$
Then $S$ has no prime objects.∎
The above result does not, however, apply with $C$ taken to be the category of all finite semigroups, since the condition that $C$ contain all semigroups $S$ with zero element equal to all three-fold products is not limited to finite $S.$ Indeed, the proof of Lemma 2.4 used semigroups $\textup{Null}(\kappa)$ in a way that required $\kappa$ to be infinite even when the given semigroup $S$ was finite. Perhaps there is some variant argument that would use only finite semigroups in such cases. Not knowing whether there is, I ask
Question 2.9.
Does the category of finite semigroups have any prime objects?
3. Preparation for some positive results
I obtained the results of the preceding section after a long grueling attempt to prove that the additive semigroup of positive integers was prime in $\mathbf{Semigp}.$ Since the argument I eventually found showing that this was not true makes strong use of action-equivalent elements, it seemed plausible that positive results might still hold in the category of semigroups with cancellation.
And, indeed, the arguments I was trying to use do work for such semigroups, and can be adapted to several weaker and variant hypotheses.
Some of the preliminary results we will develop to help us show this are true in more general contexts than those in which we will use them. In particular, though we will eventually use the next lemma with $G$ the additive group of integers, the proof of the lemma does not require $G$ to be commutative, so I will use multiplicative notation till we specialize further.
The referee indicated that that lemma was probably known. The one published result I am now aware of that implies it is [6, Lemma 3.1], where the authors say it holds by “easy computation” but give no further details.. The proof below, suggested by Benjamin Steinberg, is neater and shorter than my original version.
Recall that an ideal of a semigroup $S$ means a nonempty subset $A\subseteq S$ which is closed under left and right multiplication by elements of $S.$
Lemma 3.1.
Let $S$ be a semigroup, $A$ an ideal of $S,$ $G$ a group, and
(3.1) $\pi:A\,\to\,G$
a semigroup homomorphism. Then $\pi$ extends uniquely to a homomorphism $\hat{\pi}:S\to G.$ Indeed, for any $s\in S$ and $x\in A,$ $\hat{\pi}(s)$ can be described both as $\pi(sx)\,\pi(x)^{-1}$ and as $\pi(x)^{-1}\pi(xs).$
Proof.
Choose any $x\in A,$ and define
(3.2) $\hat{\pi}(s)~=~\pi(x)^{-1}\,\pi(x\,s\,x)~\pi(x)^{-1}$ for $s\in S.$
Since the xsxx\,s\,x factors in $A$ as $x\,s\cdot x$ and $x\cdot s\,x,$ and $\pi$ is a homomorphism on $A,$ we get corresponding factorizations of $\pi(x\,s\,x).$ Applying these to the right-hand side of (3.2), and cancelling the resulting adjacent occurrences of $\pi(x)$ and $\pi(x)^{-1},$ we get
(3.3) $\hat{\pi}(s)~=~\pi(x)^{-1}\pi(x\,s)~=~\pi(s\,x)\,\pi(x)^{-1}.$
To show that $\hat{\pi}:\,S\to G$ is a homomorphism, let $s,\,t\in S,$ put sts\,t in place of $s$ in (3.2), note that xstx=(xs)(tx)x\,s\,t\,x=(x\,s)(t\,x) with both factors in $A,$ yielding a factorization of the middle term of the right-hand side of that case of (3.2). Simplifying the result using (3.3), we get
(3.4) $\hat{\pi}(s\,t)~=~\hat{\pi}(s)\,\hat{\pi}(t)$ for all $s,\,t\in A.$
It is clear from (3.3) that $\hat{\pi}$, restricted to $A,$ gives $\pi,$ and is the unique homomorphism with this property. ∎
(The analog of the above lemma with the group $G$ replaced by a semigroup, or even a monoid, is not true. For instance, for any $n\geq 1,$ let $S$ be the additive semigroup of all positive integers, $A$ the subsemigroup of integers >n,>n, $G$ the additive monoid $A\cup\{0\},$ and $\pi:A\to G$ the inclusion map. Likewise, the analog of the lemma with the assumption that $A$ be an ideal of $S$ weakened to say that it is a left or right ideal fails. For instance, if $S$ is the free semigroup on two elements $s$ and $t,$ and we take for $A$ the left ideal generated by $t,$ then we find that $A$ is free as a semigroup on the set $\{s^{i}\,t\,|\,i\geq 0\},$ so for any nontrivial group $G$ there exist homomorphisms $\pi:A\to G$ acting arbitrarily on these generators. But such a homomorphism not satisfying the relations $\pi(s^{i}\,t)\,\pi(t)^{-1}\pi(s^{j}t)=\pi(s^{i+j}t)$ for all natural numbers $i$ and $j$ cannot be extended to $S.)$
Corollary 3.2.
Let $(S_{i})_{i\in I}$ be a family of semigroups.
For each $i\in I,$ let $S^{\prime}_{i}=S_{i}\cup\{e\},$ the monoid obtained by adjoining an identity element to $S.$ Then for any group $G,$ every semigroup homomorphism $\pi:\prod_{I}S_{i}\to G$ has a unique extension to a monoid homomorphism $\hat{\pi}:\prod_{I}S^{\prime}_{i}\to G.$
Proof.
Letting $S=\prod_{I}S^{\prime}_{i}$ and $A=\prod_{I}S_{i},$ it is easy to check that $A$ is an ideal of $S.$ Lemma 3.1 now gives the desired conclusion. ∎
Corollary 3.3.
If $S$ and $T$ are semigroups, $G$ a group, and $\pi:S\times T\to G$ a semigroup homomorphism, then there exist unique semigroup homomorphisms $\pi_{S}:S\to G$ and $\pi_{T}:T\to G$ such that the given homomorphism $\pi$ is described by
(3.5) $\pi(s,t)~=~\pi_{S}(s)~\pi_{T}(t)$ $(s\in S,\,\,t\in T).$
Moreover, $\pi_{S}(S)$ and $\pi_{T}(T)$ centralize one another in $G,$
Proof.
Apply Corollary 3.2 with $I=\{0,1\},$ $S_{0}=S,$ $S_{1}=T,$ and for $s\in S,$ $t\in T,$ define $\pi_{S}(s)=\hat{\pi}(s,e)$ and $\pi_{T}(t)=\hat{\pi}(e,t).$ From the fact that $\hat{\pi}$ respects multiplication, (3.5) and the mutual centralization assertion both follow. Uniqueness can be seen by noting that any $\pi_{S}$ and $\pi_{T}$ with the indicated properties extend to homomorphisms $S\cup\{e\}\to G$ and $T\cup\{e\}\to G,$ and calling on the uniqueness condition of Corollary 3.2. ∎
(Remark: The corresponding statement with “monoids” in place of “semigroups” is trivial – just define $\pi_{S}(s)=\pi(s,e)$ and $\pi_{T}(t)=\pi(e,t);$ and this even works with $G$ also assumed a monoid, not necessarily a group. But for $S$ and $T$ semigroups not assumed to have neutral elements $e,$ we need the above more roundabout development, including the assumption that $G$ is a group.)
[…]
Lemma 3.4.
In the context of Lemma 3.1, if $G$ is an additive subgroup of the real numbers, and $\pi:A\to G$ assumes only nonnegative values, then $\hat{\pi}:S\to G$ also assumes only nonnegative values.
Hence, likewise, in the contexts of Corollaries 3.2 and 3.3, if $G$ is an additive subgroup of the reals and $\pi$ assumes only nonnegative values, then $\hat{\pi}$, respectively $\pi_{S}$ and $\pi_{T},$ also assume only nonnegative values.
Proof.
Suppose, in the context of Lemma 3.1 with $G$ a subgroup of the real numbers, that some $s\in S$ had $\hat{\pi}(s)<0.$ Then taking any $x\in A,$ there would be a positive integer $n$ such that $\pi(s^{n}x)=n\,\hat{\pi}(s)+\pi(x)<0,$ contradicting our hypothesis on $\pi.$
Applying this in the contexts of Corollaries 3.2 and 3.3, we get the final assertion. ∎
We are ready to hone in on the class of cases we are interested in.
(3.6) For the remainder of this section, $k$ will be a fixed nonnegative integer, and $\mathbb{N}$ the additive semigroup of all nonnegative integers, so that $\mathbb{N}\!+\!k$ will denote the additive semigroup of integers $\geq k.$
We can now give necessary and sufficient conditions for a semigroup to have $\mathbb{N}\!+\!k$ as a direct factor.
Lemma 3.5.
Let $S$ be a semigroup given with a surjective semigroup homomorphism $\pi:S\,\to\,\mathbb{N}\!+\!k,$ and a set-map $\nu:S\to S$ satisfying the following three conditions:
(3.7) For all $s\in S,$ $\pi(\nu(s))\,=\,\pi(s)+1.$
(3.8) $\nu$ gives a bijection $S\to\{s\in S\,\,|\,\,\pi(s)\,\geq\,k\!+\!1\}.$
(3.9) For all $s,\,s^{\prime}\in S,$ $\nu(s)\,s^{\prime}~=~\nu(s\,s^{\prime})~=~s\,\,\nu(s^{\prime}).$
Then there exists an isomorphism of $S$ with the direct product $(\mathbb{N}\!+\!k)\times U$ of the additive semigroup $\mathbb{N}\!+\!k$ and a semigroup $U,$ such that $\pi$ corresponds to the projection to $\mathbb{N}\!+\!k,$ and $\nu$ to the map $(i,s)\mapsto(i\!+\!1,s).$
Hence, a semigroup $S$ has a direct product decomposition with $\mathbb{N}\!+\!k$ as a factor if and only if it admits maps $\pi$ and $\nu$ as above.
Proof.
Given $S,~\pi,$ and $\nu$ as above, note that (3.7) and (3.8) together imply that for every $i\geq 0,$ we have the same statements with $\nu$ replaced by $\nu^{i}$ and “+1+1” replaced by “+i+i”; and that (3.9) likewise holds with $\nu$ replaced by $\nu^{i}.$
Now define
(3.10) $U~=~\{s\in S~|~\pi(s)=k\}.$
Since $\pi$ is a semigroup homomorphism, its value on a product of two elements of $U$ will be $2k,$ hence by the $\nu^{k}\!-$version” of (3.8) we can define a binary operation on $U$:
(3.11) $u\cdot v~=~\nu^{-k}(u\,v).$
From (3.9) and the associativity of $S,$ it is easy to check that for all $u,v,w\in U,$ $(u\cdot v)\cdot w=\nu^{-2k}(u\,v\,w)=u\cdot(v\cdot w),$ so (3.11) defines a semigroup structure; and to verify that as a semigroup, $S$ is isomorphic to $(\mathbb{N}\!+\!k)\times U$ via the map
(3.12) $s~\mapsto~(\pi(s),\ \nu^{-\pi(s)+k}(s)).$
This proves the main assertion of the lemma, namely the “if” direction of the final sentence. The forward implication of that final sentence is straightforward. ∎
Below, by a direct factor of a semigroup $S$ we understand a semigroup $P$ such that $S$ has a direct product decomposition $S=~P\times Q.$ In particular, Lemma 3.5 above characterizes the semigroups having $\mathbb{N}\!+\!k$ as a direct factor.
Convention 3.6.
For the remainder of this section we shall assume that $S$ and $T$ are semigroups such that $S\times T$ has $\mathbb{N}\!+\!k$ as a direct factor. We will not write down that factorization, but simply assume we are given maps $\pi:S\times T\to\mathbb{N}\!+\!k$ and $\nu:S\times T\to S\times T$ satisfying the conditions of Lemma 3.5 with $S\times T$ in the role of $S.$
In view of Corollary 3.3 and Lemma 3.4, the map $\pi$ has the form
(3.13) $\pi(s,t)=\pi_{S}(s)+\pi_{T}(t),$ where $\pi_{S}:S\to\mathbb{N}$ and $\pi_{T}:T\to\mathbb{N}$ are semigroup homomorphisms.
We now turn to the properties of the $\nu$ of Convention 3.6. For the remainder of this section
(3.14) for $(s,t)\in S\times T,$ we shall write $\nu(s,t)=(\nu_{t}(s),\,\nu_{s}(t))\in S\times T,$ and
(3.15) $S^{\prime}$ (resp. $T^{\prime})$ will denote the subsemigroup of elements of $S$ (resp. $T)$ that can be written as products of two elements of that semigroup.
Lemma 3.7.
Assuming Convention 3.6 and the notations (3.13)-(3.15),
(3.16) for all $s\in S^{\prime},~t_{0},\,t_{1}\in T^{\prime},$ one has $\nu_{t_{0}}(s)=\nu_{t_{1}}(s),$ and their common value again lies in $S^{\prime},$ and likewise for all $s_{0},\,s_{1}\in S^{\prime},~t\in T^{\prime},$ one has $\nu_{s_{0}}(t)=\nu_{s_{1}}(t),$ and their common value again lies in $T^{\prime}.$
Hence we will make the simplification of notation
(3.17) for $s\in S^{\prime},$ we shall denote by $\nu(s)$ the common value of $\nu_{t}(s)\in S^{\prime}$ for all $t\in T^{\prime},$ and for $t\in T^{\prime},$ we shall denote by $\nu(t)$ the common value of $\nu_{s}(t)\in T^{\prime}$ for all $s\in S^{\prime},$
so that
(3.18) for all $s\in S^{\prime},~t\in T^{\prime},$ $\nu(s,t)=(\nu(s),\,\nu(t)).$
Moreover,
(3.19) for all $s,\,s^{\prime}\in S$ and $t\in T,$ $\nu_{t}(s)\,s^{\prime}~=~\nu(s\,s^{\prime})~=~s\,\nu_{t}(s^{\prime}),$ and for all $s\in S$ and $t,\,t^{\prime}\in T,$ $\nu_{s}(t)\,t^{\prime}~=~\nu(t\,t^{\prime})~=~t\,\nu_{s}(t^{\prime}).$
Proof.
For any $s,\,s^{\prime}\in S$ and $t,\,t^{\prime}\in T,$ we have, by Convention 3.6 and (3.9), $\nu(ss^{\prime},\,tt^{\prime})=\nu(s,t)\,(s^{\prime},t^{\prime}).$ Applying (3.14) to the first factor on the right-hand side, this gives $\nu(ss^{\prime},\,tt^{\prime})=(\nu_{t}(s)\,s^{\prime},\,\nu_{s}(t^{\prime})\,t^{\prime}).$ Taking first components, we have
(3.20) $\nu_{tt^{\prime}}(s\,s^{\prime})~=~\nu_{t}(s)\,s^{\prime}.$
Note that the right-hand side does not depend on $t^{\prime},$ hence neither does the left-hand side; and by the right-left dual argument, it likewise does not depend on $t,$ so we get the same value on replacing the subscript $t\,t^{\prime}$ by any element of $T^{\prime}.$ Here the argument of $\nu,$ a product of two arbitrary elements of $S,$ can be any element of $S^{\prime}.$ This proves the first line of (3.16), allowing us to adopt the notation in the first line of (3.17). The first equality of the first line of (3.19) follows; and by interchanging right and left multiplication, and the roles of $S$ and $T,$ as appropriate, we get the remaining equalities of (3.16) and (3.19). ∎
At this point, we cannot assert (3.16) without the restriction of the arguments to $S^{\prime}$ and $T^{\prime}.$ But we can deduce a strong result on the behavior of $\pi_{S}$ and $\pi_{T}$ on arbitrary elements:
Lemma 3.8.
For $S,$ $T,$ $\pi$ and $\nu$ as in Convention 3.6, we have either
(3.21) for all $s\in S,~t\in T,$ $\pi_{S}(\nu_{t}(s))~=~\pi_{S}(s)+1\,$ and  $\pi_{T}(\nu_{s}(t))~=~\pi_{T}(t),$ or for all $s\in S,~t\in T,$ $\pi_{S}(\nu_{t}(s))~=~\pi_{S}(s)\,$ and  $\pi_{T}(\nu_{s}(t))~=~\pi_{T}(t)+1.$
Proof.
Let us fix arbitrary elements
(3.22) $s_{0}\in S^{\prime},$ $t_{0}\in T^{\prime},$
and let $m$ and $n$ be the integers such that
(3.23) $\pi_{S}(\nu(s_{0}))~=~\pi_{S}(s_{0})+m$  and  $\pi_{T}(\nu(t_{0}))~=~\pi_{T}(t_{0})+n.$
Applying (3.7) (with $S\times T$ in the role of $S)$ and (3.13) to $(s_{0},\,t_{0})$ we see that
(3.24) m+n= 1.m+n\,=\,1.
Now note that for any $s\in S,~t\in T,$ we have the following equalities (using, at the first and third steps, the fact that $\pi_{S}$ is a homomorphism, at the second, (3.19), and at the last, (3.23)):
(3.25) $\pi_{S}(\nu_{t}(s))+\pi_{S}(s_{0})=\pi_{S}(\nu_{t}(s)\,s_{0})=\pi_{S}(s\,\nu(s_{0}))=\pi_{S}(s)+\pi_{S}(\nu(s_{0}))=\pi_{S}(s)+\pi_{S}(s_{0})+m.$
Cancelling $\pi_{S}(s_{0})$ at the beginning and end, we get
(3.26) $\pi_{S}(\nu_{t}(s))~=~\pi_{S}(s)+m$ for all $s\in S,$ $t\in T.$
Reversing the roles of $S$ and $T,$ we likewise have
(3.27) $\pi_{T}(\nu_{s}(t))~=~\pi_{T}(t)+n$ for all $s\in S,$ $t\in T.$
Now by (3.13), $\pi_{S}$ assumes only nonnegative values, so we can choose an $s\in S$ minimizing $\pi_{S}(s),$ and applying (3.26) to this $s,$ we see that $m$ must be nonnegative; and the analogous argument shows the same for $n.$ From (3.24), it follows that one of $m$ and $n$ is 0 and the other $1.$ Substituting into (3.26) and (3.27) we get (3.21). ∎
Corollary 3.9.
If the first alternative of (3.21) holds, then the function $\pi_{T}$ is identically 0 and $\pi_{S}(S)=\mathbb{N}\!+\!k,$ while if the second holds, $\pi_{S}$ is identically 0 and $\pi_{T}(T)=\mathbb{N}\!+\!k.$
Proof.
Assume without loss of generality that the first alternative of (3.21) holds, and suppose $\pi_{T}$ were not identically zero. Then it would assume some positive value, and being a semigroup homomorphism, would assume arbitrarily large values.
Let $s\in S$ be an element at which $\pi_{S}$ takes on its least value. By the above observation, we can choose $t\in T$ such that the value of $\pi(s,\,t)=\pi_{S}(s)+\pi_{T}(t)$ is $\geq k\!+\!1.$
Then by (3.8) applied to $S\times T,$ this implies that for some $(s^{\prime},t^{\prime})\in S\times T,$ $(s,t)=\nu(s^{\prime},t^{\prime})=(\nu_{t^{\prime}}(s^{\prime}),\nu_{s^{\prime}}(t^{\prime}))$ By the case of (3.21) we are assuming, this gives $\pi_{S}(s)=\pi_{S}(\nu_{t^{\prime}}(s^{\prime}))=\pi_{S}(s^{\prime})+1,$ so $\pi_{S}(s^{\prime})<\pi_{S}(s),$ contradicting our choice of $s.$
So $\pi_{T}$ is indeed identically zero. Substituting this into (3.13) and observing that $\pi(S\times T)=\mathbb{N}\!+\!k$ by Convention 3.6 and the conditions of Lemma 3.5 it refers to, we conclude that $\pi_{S}(S)=\mathbb{N}\!+\!k,$ as claimed. ∎
4. Primeness results, at last
We are now ready to note some subcategories of $\mathbf{Semigp}$ in which we can use the above results to prove some or all of the semigroups $\mathbb{N}\!+\!k$ prime. These categories cannot, of course, contain the semigroups $\textup{Null}(\kappa)$ that were used to prove non -primeness in §2. Here are some restrictions that suffice.
Definition 4.1.
We shall call a semigroup $S$
(i) cancellative if it satisfies
(4.1) for all $s\neq s^{\prime}\in S$ and $s^{\prime\prime}\in S,$ one has $s\,s^{\prime\prime}\neq s^{\prime}\,s^{\prime\prime},$ and
(4.2) for all $s\neq s^{\prime}\in S$ and $s^{\prime\prime}\in S,$ one has $s^{\prime\prime}\,s\neq s^{\prime\prime}\,s^{\prime},$
(ii) right cancellative if it merely satisfies (4.1), respectively, left cancellative if it merely satisfies (4.2),
and,
(iii) weakly cancellative if it satisfies
(4.3) for all $s\neq s^{\prime}\in S$ there exists an $s^{\prime\prime}\in S$ such that one of the inequalities $s\,s^{\prime\prime}\neq s^{\prime}\,s^{\prime\prime}$ or $s^{\prime\prime}\,s\neq s^{\prime\prime}\,s^{\prime}$ holds
(i.e. if, in the language of Definition 2.2(ii), no two distinct elements of $S$ are action equivalent).
It is easy to verify
Lemma 4.2.
For each of conditions (4.1), (4.2), (4.3), a direct product semigroup $S\times T$ satisfies the condition in question if and only if both $S$ and $T$ do.∎
We can now prove
Theorem 4.3.
For all nonnegative integers $k,$ the semigroup $\mathbb{N}\!+\!k$ is prime in the category of cancellative semigroups, and, more generally, in the categories of right cancellative and left cancellative semigroups, and still more generally, in the category of weakly cancellative semigroups.
Proof.
We shall prove $\mathbb{N}\!+\!k$ prime in the category of weakly cancellative semigroups. In view of Lemma 4.2, the other categories mentioned are closed in that one under pairwise direct products and direct factors, hence primeness of $\mathbb{N}\!+\!k$ in that category will imply primeness in the others.
So suppose $S$ and $T$ are weakly cancellative semigroups such that $\mathbb{N}\!+\!k$ is a direct factor of $S\times T.$ As shown in the preceding section, this is equivalent to the existence of maps $\pi:S\times T\to\mathbb{N}\!+\!k$ and $\nu:S\times T\to S\times T$ satisfying (3.7)-(3.9). As in (3.14), for $(s,t)\in S\times T,$ we shall write $\nu(s,t)=(\nu_{t}(s),\,\nu_{s}(t)).$
We claim that under our present hypotheses, $\nu_{t}(s)$ must in fact be a function of $s$ alone, and $\nu_{s}(t)$ a function of $t$ alone. By symmetry, it suffices to prove the former statement.
So suppose that for some $s\in S,$ $t,\,t^{\prime}\in T$ we had $\nu_{t}(s)\neq\nu_{t^{\prime}}(s).$ We now invoke weak cancellativity. By left-right symmetry with respect to order of operations, we may assume without loss of generality that there is some $s^{\prime}\in S$ such that $\nu_{t}(s)\,s^{\prime}\neq\nu_{t^{\prime}}(s)\,s^{\prime}.$ But by (3.19), both sides equal $\nu(s\,s^{\prime}),$ a contradiction.
Hence the operators $\nu_{t}\!:S\to S$ $(t\in T)$ are all equal, and likewise the operators $\nu_{s}\!:T\to T$ $(s\in S).$ Writing the common values of each as $\nu:S\to S$ and $\nu:T\to T,$ (3.14) becomes
(4.4) for all $s\in S,$ $t\in T,$ $\nu(s,t)=(\nu(s),\,\nu(t)).$
Now by Lemma 3.8, one of the functions $\pi_{S},$ $\pi_{T}$ adds $1$ whenever $\nu$ is applied to its argument; without loss of generality, let us assume that is $\pi_{S}.$ Then by Corollary 3.9, $\pi_{S}$ takes on all the values in $\mathbb{N}\!+\!k$ and $\pi_{T}$ is identically $0.$
From the fact that the maps $\pi$ and $\nu$ on $S\times T$ satisfy the conditions (3.7)-(3.9) of Lemma 3.5 (with $S\times T$ in the role of $S),$ we now see that $\pi_{S}:S\to\mathbb{N}\!+\!k$ and $\nu:S\to S$ satisfy those same conditions, hence, by that lemma, $\mathbb{N}\!+\!k$ is a direct factor in $S.$ (If, rather, $\pi_{T}$ is the function that, as indicated in the preceding paragraph, adds $1$ when $\nu$ is applied to its argument, then $T$ has $\mathbb{N}\!+\!k$ as a direct factor.) ∎
Our other result is easier to prove, but applies only to $\mathbb{N},$ since for $k>0,$ $\mathbb{N}\!+\!k$ does not belong to the categories in question.
Theorem 4.4.
The object $\mathbb{N}$ is prime in the full subcategory of $\mathbf{Semigp}$ whose objects are the monoids, and, more generally, in the full subcategory of $\mathbf{Semigp}$ whose objects are the semigroups in which every element is a product.
Proof.
In the latter category, Lemma 3.7 shows that all the functions $\nu_{t}:S\to S$ are equal, and likewise the functions $\nu_{s}:T\to T.$ This gives us (4.4), and we complete the proof as for Theorem 4.3. ∎
The monoid case of Theorem 4.4 is also implied by the “weakly cancellative” case of Theorem 4.3, since multiplication by the identity element is cancellable.
We can easily go from the above result about monoids as a full subcategory of $\mathbf{Semigp}$ to one about the category of monoids and monoid homomorphisms:
Proposition 4.5.
A monoid $S$ which is prime in the full subcategory of $\mathbf{Semigp}$ whose objects are the monoids is also prime in the category of monoids (i.e., the subcategory thereof where morphisms are required to respect identity elements). In particular, $\mathbb{N}$ is prime in that category.
Sketch of proof.
The direct product $S\times T$ of two objects in the category of monoids clearly has the same semigroup structure as their direct product in the category of semigroups. Also, if $S$ and $T$ are semigroups such that $S\times T$ is a monoid, it is easy to verify that $S$ and $T$ must be monoids. Hence a monoid $S$ is a direct factor of a monoid $U$ in the category of semigroups if and only if it is a direct factor of $U$ in the category of monoids.
From these observations, it is easy to deduce the first assertion of this Proposition, and apply it to the monoid $\mathbb{N}.$
However, the referee has pointed out that the above property of $\mathbb{N}$ is a case of a much more general known result.
The context of that result concerns categories of algebras whose operations include a binary operation denoted ++ (but not required to be commutative or even associative) and a zeroary operation $0,$ such that $\{0\}$ is a subalgebra, and the identity $0+x=x=x+0$ holds [9, start of §5.4]. The result also involves the “center” of such an algebra, a complicated concept developed in [9, §5.5]; but all we will need to know about it below is that for all elements $b$ in the center, there exists an element $c$ such that $b+c=0$ [9, Definitions 5.11 and 5.12]. (Warning: the word “center” is used in different, though overlapping ways in other chapters of [9].) The result is
(4.5) [9, Exercise 5.15 (10), p. 300] In the category of all algebras of a given type with 0 and ++ as above, every nontrivial object with finite center that is not a direct product of two nontrivial objects is prime.
Note that primeness in such a category implies primeness in any subvariety, and that a monoid with no nontrivial invertible elements (or more generally, with only finitely many) necessarily satisfies the “finite center” condition of (4.5).
Some examples in the variety of monoids that (4.5) shows to be prime include: (i) all nontrivial submonoids of $\mathbb{N},$ (ii) all submonoids $S$ of $\mathbb{N}\times\mathbb{N}$ such that the submonoids $S\cap(\mathbb{N}\times\{0\})$ and $S\cap(\{0\}\times\mathbb{N})$ are both nontrivial, and $S$ is strictly larger than their sum (e.g., $\{(a,\,b)\in\mathbb{N}\times\mathbb{N}\ |\ a\equiv b\ (\mathrm{mod}~n)\}$ for any $n>1;$ or $\mathbb{N}\times\mathbb{N}\setminus\{(1,0)\}),$ (iii) all submonoids $S$ of $\mathbb{N}\times\mathbb{N}$ such that $S\cap(\mathbb{N}\times\{0\})=\{(0,0)\},$ while $S\cap(\mathbb{N}\times\{1\})$ and $S\cap(\{0\}\times\mathbb{N})$ each have more than one element, and (iv) all free monoids on more than one (possibly infinitely many) generators. (That none of these examples is a nontrivial directly product takes a bit of thought in case (ii) and more in case (iii). I leave these as exercises for the reader.)
If one is interested in monoids including nontrivial invertible elements, it is worth noting that the concept of “center” used in (4.5) has the property that every central element of an algebra is central in the more familiar sense that it commutes, under +,+, with every element of the algebra [9, p. 296, sentence containing first display]. This leads to still larger classes of objects with finite centers in the categories of monoids and of groups.
On the other hand, we shall see in Example 6.2 below that the additive group of integers is not prime (in the language of that section, “not Tarski-prime”) in the category of groups, from which it easily follows that it is also non-prime in the categories of semigroups and monoids considered above.
5. Two other questions from [9]
The question answered in §2 above is the second of four questions listed in [9, middle of p. 263]. It was noted in [14] that the last of those questions, “Is $\mathbb{Z}$ prime in the class of all Abelian groups?”, has an easy affirmative answer. So far as I know, the other two remain open.
One of them falls under the general theme of this note. In it, an idempotent semigroup means one satisfying the identity xx=x.x\,x=x. There is considerable literature on such semigroups, e.g. [4, §4.4] and [2].
Question 5.1.
[9, p. 263] Is there any prime in the class of all idempotent semigroups?
The other question has a less obvious connection with this note. In it, an algebra “of type <1,1><\!1,1\!>” means an algebra having precisely two operations, both unary.
Question 5.2.
[9, p. 263] Is there any prime in the class of all finite algebras of type <1,1><\!1,1\!> (or in any type that has more than one operation of arity $\geq 1)$?
An algebra of type <1,1><\!1,1\!> can be thought of as a set given with an action on it of the free monoid on two generators; so in that way the question has a somewhat similar flavor to those studied in the preceding sections.
6. Some results and examples for groups, comparing
Tarski-primeness and the two sorts of Rhodes-primeness
We mentioned in §1 that two concepts called primeness, somewhat related to but distinct from Tarski’s, are studied by semigroup theorists. In this section, we examine the relationships among these three sorts of primeness when applied to groups. We give them distinct names in Convention 6.1(iii)-(v) below.
This section can be read independent of the preceding material (ignoring a couple of brief comments about the contrast between Example 6.2 below and Theorems 4.3 and 4.4 above, and, much later, some of the remarks between Questions 6.13 and 6.14).
Convention 6.1.
In this section, for objects $G,$ $H,$ etc. of the category of all groups, or the category of finite groups,
(i)     we shall call $H$ a direct factor of $G$ if $G$ is isomorphic to a direct product $H\times H^{\prime},$
(ii)   we shall call $H$ a subquotient of $G$ if $H$ is isomorphic to a homomorphic image of a subgroup of $G.$
Moreover, for $G$ nontrivial,
(iii) we shall call $G$ Tarski-prime if, whenever $G$ is a direct factor of a group in the category in question, and the latter group can also be written as a direct product $G_{0}\times G_{1},$ then $G$ is a direct factor of one of $G_{0}$ or $G_{1},$
(iv)  we shall call $G$ Rhodes-prime with respect to direct products if, whenever $G$ is a subquotient of a group in the category in question, and the latter group can be written as a direct product $G_{0}\times G_{1},$ then $G$ is a subquotient of one of $G_{0}$ or $G_{1}$ (condition introduced at [11, p. 482, line 6]); and similarly,
(v) we shall call $G$ Rhodes-prime with respect to semidirect products if, whenever $G$ is a subquotient of a group in the category in question, and the latter group can be written as a semidirect product $G_{0}\rtimes\,G_{1},$ then $G$ is a subquotient of one of $G_{0}$ or $G_{1}$ (condition introduced at [11, p. 228, paragraph before Lemma 4.1.31]).
(What we name a subquotient of $G$ in (ii) above was earlier also called a factor of $G,$ e.g., [10, p. 1 et seq.], is called a divisor of $G$ in [11, Def. 1.2.25, p. 32], and is now often called a section of $G$ by group theorists. But “subquotient” is also fairly common.)
We will also follow the common conventions of generally writing $e$ for the identity elements of groups, though for particular groups we may use notation appropriate for them, e.g., 0 for the identity element of the additive group $\mathbb{Z};$ and of calling a group $G$ nontrivial if it has more than one element. (We use “nontrivial” in that sense in the line introducing (iii)-(v) above. In [11] the trivial semigroup is counted as prime with respect to direct and semidirect products, but here we have excluded that group to make the comparison between the Rhodes-primeness conditions and that of Tarski-primeness more straightforward.)
In view of Theorems 4.3 and 4.4 of the preceding section, one might expect the additive group $\mathbb{Z}$ to be Tarski-prime in the category of all groups. But a key tool in proving those results was the fact that if two members of $\mathbb{N}$ satisfy $m+n=1$ (3.24), one of them has to be $0.$ The failure of that statement in $\mathbb{Z}$ underlies the difference in its behavior.
Example 6.2.
In the category of groups, $\mathbb{Z}$ is not Tarski-prime, but is Rhodes-prime with respect to semidirect and direct products.
Proof.
To construct a counterexample to Tarski-primeness, let us choose any
(6.1) prime numbers $p$ and $q,$ and nonunit divisors a|p1a\,|\,p\,{-}1 and b|q1,b\,|\,q\,{-}1, such that $a$ and $b$ are relatively prime.
(E.g., $p=3,$ $q=7,$ $a=2,$ $b=3.$ Or one could take both $p$ and $q$ to be $7,$ again with $a=2$ and $b=3.)$
Recalling that the multiplicative groups of the rings $\mathbb{Z}_{p}$ and $\mathbb{Z}_{q}$ are cyclic of orders p1p-1 and q1,q-1, we can
(6.2) let $G=(\mathbb{Z}_{p}\rtimes\mathbb{Z})\times(\mathbb{Z}_{q}\rtimes\mathbb{Z})$ where, in the first factor, a generator of $\mathbb{Z}$ acts on $\mathbb{Z}_{p}$ by multiplication by a unit of the ring $\mathbb{Z}_{p}$ having multiplicative order $a,$ and in the second, such a generator acts on $\mathbb{Z}_{q}$ by multiplication by a unit of the ring $\mathbb{Z}_{q}$ having multiplicative order $b.$
Note that neither of the groups $\mathbb{Z}_{p}\rtimes\mathbb{Z}$ and $\mathbb{Z}_{q}\rtimes\mathbb{Z}$ has a direct product decomposition with a factor isomorphic to $\mathbb{Z}.$ E.g., if $\mathbb{Z}_{p}\rtimes\mathbb{Z}$ did, then a generator of that factor would be a central element of infinite order, hence would have, under the given description of the group, the form (0,ma)(0,\,ma) $(m\neq 0).$ But that would make it an a\!a\!-th power in the whole group, while looking at the $\mathbb{Z}\!$-component in such a direct product decomposition, we see that a generator thereof can’t be an a\!a\!-th power for any $a>1.$ (The groups $\mathbb{Z}_{p}\rtimes\mathbb{Z}$ and $\mathbb{Z}_{q}\rtimes\mathbb{Z}$ are in fact directly indecomposable, but the above argument gives as much as we need.)
On the other hand, suppose we rewrite $G$ as $(\mathbb{Z}_{p}\times\mathbb{Z}_{q})\rtimes(\mathbb{Z}\times\mathbb{Z}),$ using the action of $\mathbb{Z}\times\mathbb{Z}$ under which the first factor $\mathbb{Z}$ affects only the first factor of $\mathbb{Z}_{p}\times\mathbb{Z}_{q},$ and the second affects only the second.
Since $a$ and $b$ are relatively prime, the subgroup of $\mathbb{Z}\times\mathbb{Z}$ generated by the element $(a,b)$ is a direct factor thereof. Explicitly, we can find integers $c$ and $d$ such that $bc+ad=1,$ and use the right action of the invertible matrix $(\begin{smallmatrix}\ b&d\\ \!\!-a&c\end{smallmatrix})$ to get an isomorphism $\mathbb{Z}\times\mathbb{Z}\cong\mathbb{Z}\times\mathbb{Z}$ such that the image of $(a,b)$ generates the second factor. (That matrix or its inverse necessarily has some negative entries, which can be thought of as the way the fact that $\mathbb{Z}$ includes negative integers comes in.) We see from (6.2) that under the action on $\mathbb{Z}_{p}\times\mathbb{Z}_{q},$ that second factor acts trivially.
Thus, $G$ can be written as a semidirect product $G=(\mathbb{Z}_{p}\times\mathbb{Z}_{q})\rtimes(\mathbb{Z}\times\mathbb{Z})$ where the second $\mathbb{Z}$ acts trivially, and all the action comes from the first. Hence we can rewrite this group as a direct product $G=((\mathbb{Z}_{p}\times\mathbb{Z}_{q})\rtimes\mathbb{Z})\times\mathbb{Z}.$ So the group $\mathbb{Z}$ is isomorphic to a direct factor of the product group (6.2), though it is not a direct factor in either of the given factors. So $\mathbb{Z}$ is not Tarski-prime.
As to the Rhodes-primeness conditions, it is clear that $\mathbb{Z}$ is a subquotient of a group $G$ if and only if $G$ has an element of infinite order. It is also easy to see that a direct or semidirect product $G_{0}\times G_{1}$ or $G_{0}\rtimes G_{1}$ has an element if infinite order if and only if $G_{0}$ or $G_{1}$ does. So $\mathbb{Z}$ satisfies both versions of Rhodes-primeness. ∎
For finite groups, the relationship between these primeness conditions is quite different. A key result is the Krull-Schmidt Theorem [5, Theorem 3.8, p. 86], which says that every group with ACC and DCC on normal subgroups has a factorization as a direct product of finitely many nontrivial groups which are not themselves nontrivial direct products, and that the factors in such a decomposition are unique up to rearrangement and isomorphism. Thus, every group with those chain conditions which is not a nontrivial direct product is Tarski-prime in the category of such groups. In particular, every nontrivial finite group which is not a nontrivial direct product is Tarski-prime in the category of finite groups. (For a result about finite algebras of much more general sorts than groups, of which this is a special case, see (4.5) above, together with the paragraph that precedes that display, and the second paragraph following it.)
Recall that if a group $G$ has a unique minimal nontrivial normal subgroup $M,$ then $G$ is called monolithic, and $M$ is called the monolith of $G$ [10, sentence beginning at bottom of p. 146]. We can now prove
Lemma 6.3.
In the category of finite groups, for $G$ a nontrivial group we have the implications
(6.3) $G$ is Rhodes-prime with respect to semidirect products $\implies$ $G$ is Rhodes-prime with respect to direct products $\implies$ $G$ is monolithic $\implies$ $G$ is Tarski-prime.
Proof.
The first implication is clear from the definitions, since direct products are a special case of semidirect products.
We shall prove the remaining two implications by contradiction. For the middle implication, suppose $G$ has two minimal nontrivial normal subgroups $N\neq N^{\prime}.$ Then $N\cap N^{\prime}=\{e\},$ so the natural map $G\to(G/N)\times(G/N^{\prime})$ is an embedding. Thus $G$ is isomorphic to a subgroup of the direct product $(G/N)\times(G/N^{\prime}),$ making it a subquotient thereof; but it is not a subquotient of either $G/N$ or $G/N^{\prime},$ since these have smaller orders than $G.$ So $G$ is not Rhodes-prime with respect to direct products.
For the final implication, note that by the Krull-Schmidt Theorem, every non-Tarski-prime nontrivial group $G$ is a direct product $G_{0}\times G_{1}$ of nontrivial groups, and if we choose minimal nontrivial normal subgroups $N_{0}\subseteq G_{0}$ and $N_{1}\subseteq G_{1},$ then $N_{0}\times\{e\}$ and $\{e\}\times N_{1}$ are distinct minimal nontrivial normal subgroups of $G,$ so $G$ is not monolithic. ∎
Now for some positive results on when finite groups are Rhodes-prime, in each of the two senses; though our first result will not be limited to finite groups:
Proposition 6.4.
Every nontrivial simple group $G$ (commutative or noncommutative) is Rhodes-prime with respect to semidirect products in the category of all groups.
Hence every nontrivial finite simple group $G$ is Rhodes-prime with respect to semidirect products in the category of finite groups.
Proof.
Let $G$ be a nontrivial simple group, and suppose $G$ is a subquotient of a semidirect product $G_{0}\rtimes\,G_{1}.$ That is, suppose that that we can write it as $f(H)$ for some subgroup $H<G_{0}\rtimes\,G_{1},$ and some surjective homomorphism $f:H\to G.$
Since $G_{0}\rtimes\{e\}$ is a normal subgroup of $G_{0}\rtimes G_{1},$ $H\cap(G_{0}\rtimes\{e\})$ will be a normal subgroup of $H,$ so its image in $G$ under $f$ will be normal; so since $G$ is simple, that image will either be $G$ or $\{e\}.$
If $f(H\cap(G_{0}\rtimes\{e\}))=G,$ that makes $G$ a homomorphic image of a subgroup of $G_{0},$ i.e., a subquotient of $G_{0}.$
On the other hand, if $f(H\cap(G_{0}\rtimes\{e\}))=\{e\},$ we see that the image in $G$ of an element $(g_{0},g_{1})\in H$ depends only on $g_{1},$ and easily deduce that $G$ is a homomorphic image of the subgroup of $G_{1}$ consisting elements occurring as second components of members of $H;$ so $G$ is a subquotient of $G_{1}.$
This proves the Rhodes-primeness of $G$ with respect to semidirect products. If $G$ is finite, having that property in the category of all groups clearly implies that it has the same property in the category of finite groups. ∎
As noted in Example 6.2, the converse to Proposition 6.4 fails in the category of all groups: the non-simple group $\mathbb{Z}$ is Rhodes-prime with respect to semidirect products. But Alexander Olshanskiy (personal communication) has provided a proof that that converse does hold in the category of finite groups. Here is a simplified version of his proof.
Proposition 6.5 (A. Olshanskiy).
If a nontrivial finite group $G$ is Rhodes-prime with respect to semidirect products in the category of finite groups, then $G$ is simple.
Proof.
Let $N$ be a minimal nontrivial normal subgroup of $G.$ (By Lemma 6.3 there is a unique such $N,$ but we will not need to call on that fact.)
Being an extension of $N$ by $G/N,$ the group $G$ can be written as a subgroup of the wreath product $N\wr G/N,$ that is, of the semidirect product $N^{|G/N|}\rtimes G/N,$ where $|G/N|$ denotes the underlying set of $G/N,$ and $G/N$ acts on that direct power of $N$ by permutation of the factors (Theorem of Kaloujnine and Krasner, [10, Theorem 22.21, p. 46]). Hence by Rhodes-primeness with respect to semidirect products, $G$ must be a subquotient of $G/N$ or of $N^{|G/N|}.$ The former is impossible because $G/N$ has smaller order than $G,$ so $G$ must be a subquotient of $N^{|G/N|}.$
Though $N$ is a minimal nontrivial normal subgroup of $G,$ it need not be simple, but it will be characteristically simple, i.e., it will have no proper nontrivial subgroup invariant under all its automorphisms (since these include conjugation by all members of $G).$ Hence by [12, 3.3.15, p. 87, and sentence at end of proof, p. 88], $N$ is a direct product of (mutually isomorphic) simple groups. Hence $N^{|G/N|}$ is a direct product of those same simple groups, with possibly more repetitions.
Thus $G,$ being Rhodes-prime with respect to semidirect products, and hence in particular, with respect to direct products, must be a homomorphic image of a subgroup one of those simple groups, $S.$ (Our definition of that Rhodes-primeness condition only refers to pairwise products, but by induction, it extends to arbitrary finite products.) But since $S$ is a direct factor of the subgroup $N$ of $G,$ its order is less than or equal to that of $G,$ so to be a homomorphic image of a subgroup of $S,$ the group $G$ must, in fact, be isomorphic to $S,$ hence, as desired, simple. ∎
The next result gives some sufficient conditions for $G$ to be Rhodes-prime with respect to direct products. In view of the second implication of (6.3), $G$ must be monolithic. Given this, we find that some quite varied additional conditions imply the desired Rhodes-primeness. (The referee points out that case (i) below is [1, Theorem 10.1] applied to the variety of groups.)
Proposition 6.6.
Let $G$ be a monolithic finite group, with monolith $M.$
Then $G$ is Rhodes-prime with respect to direct products in the category of finite groups if any of the following conditions holds:
(i) $M$ is noncommutative, or
(ii) $G$ is a semidirect product $M\rtimes K$ of $M$ with a subgroup $K<G,$ or
(iii) $G$ is a cyclic group of prime-power order, $\mathbb{Z}_{p^{n}}.$
Proof.
Given a finite monolithic group $G$ that is not Rhodes-prime with respect to direct products, we shall show that neither (i) nor (ii) can hold. The proof that (iii) implies Rhodes-primeness will be more straightforward.
To get the first two results, suppose $G$ is not Rhodes-prime with respect to direct products, i.e., can be written as a homomorphic image $f(H)$ of a subgroup $H$ of a direct product $G_{0}\times G_{1}$ of finite groups, such that $G$ is not a subquotient of $G_{0}$ or of $G_{1}.$ We can clearly replace $G_{0}$ and $G_{1}$ by the subgroups given by the projections of $H$ onto those two factors, so assume that those projections are surjective; i.e., that $H$ is a subdirect product of $G_{0}$ and $G_{1}.$ Let us further assume that, for the given group $G,$ the groups $G_{0},$ $G_{1},$ and $H$ are chosen so as to minimize the order of $H.$
Now let
(6.4) $H_{0}=\{h\in G_{0}\,|\,(h,e)\in H\}$  and   $H_{1}=\{h\in G_{1}\,|\,(e,h)\in H\}.$
If $H_{0}$ were trivial, then no two elements of $H$ would fall together under the projection $G_{0}\times G_{1}\to G_{1},$ so $G_{1}$ would be isomorphic to $H,$ so as $G$ is a homomorphic image of $H,$ this would contradict the assumption that $G$ was not a subquotient of $G_{1}.$ So $H_{0}$ is nontrivial; and similarly $H_{1}.$ From our assumption that $H$ projects surjectively to each of $G_{0}$ and $G_{1}$ it is also easy to see that $H_{0}$ is normal in $G_{0},$ and $H_{1}$ in $G_{1}.$ (E.g., given $h\in H_{0}$ and $g_{0}\in G_{0},$ we can find $g_{1}\in G_{1}$ such that $(g_{0},g_{1})\in H;$ and conjugating $(h,e)$ by $(g_{0},g_{1}),$ we effectively conjugate $h$ by $g_{0}.)$ Now if $\mathrm{ker}(f)$ had nontrivial intersection with the subgroup $H_{0}\times\{e\}$ of $H,$ then since $H_{0}\times\{e\}$ and $\mathrm{ker}(f)$ are both normal in $H,$ their intersection would be normal, and dividing $G_{0}$ by the projection of this intersection, we could decrease the order of $H.$ This, and the corresponding observation for $\{e\}\times H_{1},$ give
(6.5) $\mathrm{ker}(f)$ has trivial intersections with $H_{0}\times\{e\}$ and with $\{e\}\times H_{1}.$
Now let $M_{0}\times\{e\}$ be a nontrivial subgroup of $H_{0}\times\{e\}$ minimal for being normal in $H.$ Since $M_{0}\subseteq H_{0},$ (6.5) tells us that $M_{0}\times\{e\}$ maps one-to-one into $G,$ hence its image is a minimal normal subgroup of $G;$ hence that image is $M.$ Combining with the corresponding observation about a minimal normal subgroup $\{e\}\times M_{1}$ of $\{e\}\times H_{1},$ we get
(6.6) $H_{0}\times\{e\}$ is monolithic with monolith $M_{0}\times\{e\},$ and $\{e\}\times H_{1}$ is monolithic with monolith $\{e\}\times M_{1},$ and both of these map isomorphically to $M$ under $f.$
But $M_{0}\times\{e\}$ and $\{e\}\times M_{1}$ centralize one another in $H,$ so $M$ must be self-centralizing in $G,$ i.e., commutative, giving the desired contradiction to (i).
To get the next assertion, suppose as in (ii) that $G$ is a semidirect product $M\rtimes K$ for some $K<G.$ Since $K$ has trivial intersection with the monolith $M$ of $G,$
(6.7) $K$ contains no nontrivial normal subgroup of $G.$
On the other hand, since $M$ is invariant under conjugation by members of $K,$ the centralizer of $M$ in $K$ must also be invariant under conjugation by members of $K;$ and by definition that centralizer will be invariant under conjugation by members of $M;$ hence it is invariant under conjugation by all members of $MK=G,$ i.e., it is normal in $G.$ So by (6.7),
(6.8) The centralizer of $M$ in $K$ is trivial.
Returning to what we proved earlier about $H,$ note that if $f^{-1}(K)<H$ had nontrivial intersection with $H_{0},$ then this would centralize $H_{1},$ hence its image in $G$ would be a subgroup of $K,$ nontrivial by (6.5), that centralized the image of $H_{1},$ which we saw in (6.6) contains $M,$ contradicting (6.8). So $f^{-1}(K)$ has trivial intersection with $H_{0};$ and similarly with $H_{1}.$ In view of (6.4), this forces $f^{-1}(K)$ to be the graph of an isomorphism between subgroups $K_{0}<G_{0}$ and $K_{1}<G_{1},$ each isomorphic to $K,$ and we see that these act on $M_{0}$ and $M_{1}$ as $K$ acts on $M.$ Hence the subgroup $M_{0}\,K_{0}<G_{0}$ (and likewise $M_{1}\,K_{1}<G_{1})$ is isomorphic to $MK=G,$ again contradicting our assumption that $G$ is not a subquotient of $G_{0}$ or $G_{1}.$ This gives the desired contradiction to (ii) for $G$ not Rhodes-prime with respect to direct products.
To prove Rhodes-primeness with respect to direct products in case (iii), suppose a subgroup $H$ of a finite group $G_{0}\times G_{1}$ can be mapped surjectively to $G=\mathbb{Z}_{p^{n}}.$ Then an element mapping to a generator of $G$ must have order divisible by $p^{n}.$ Since the order of an element of $G_{0}\times G_{1}$ is the least common multiple of the orders of its components, one of those components must have order divisible by $p^{n},$ hence some power of that element will have order exactly $p^{n},$ hence the subgroup of $G_{0}$ or $G_{1}$ that it generates will be a subquotient of $G_{0}$ or $G_{1}$ isomorphic to $\mathbb{Z}_{p^{n}}=G.$
(I obtained case (ii) of the above result in 2014, answering a question posed by John Rhodes, personal correspondence. That result is given in weakened form in [8, Theorem 4.20, p. 1275]. For the meaning of “$G$ is ji” in that statement, see the last paragraph of [8, p. 1252], in particular, the display.)
Let us note some examples of the distinction between Rhodes-primeness with respect to semidirect and direct products.
Example 6.7.
In the category of finite groups, the following groups are Rhodes-prime with respect to direct products but, not being simple, are not Rhodes-prime with respect to semidirect products.
(i)  The permutation groups $S_{n}$ for all $n\geq 3.$
(ii)  All semidirect products $\mathbb{Z}_{p}\rtimes A$ where $p$ is a prime and $A$ a nontrivial subgroup of $\mathrm{Aut}(\mathbb{Z}_{p})\cong\mathbb{Z}_{p-1}.$
(iii) All groups $\mathbb{Z}_{p^{n}}$ for $p$ prime and $n>1.$
Proof.
To show $S_{n}$ Rhodes-prime with respect to direct products, we use Proposition 6.6(ii). For all cases except $n=4,$ $S_{n}$ is monolithic with monolith $A_{n},$ and is a semidirect product of $A_{n}$ with the order-2\!2\! subgroup generated by any transposition. The group $S_{4}$ has a different monolith, the Klein four-group $V,$ consisting of the even permutations of exponent $2,$ and if we write $S_{3}$ for the subgroup of elements of $S_{4}$ fixing some one of the four elements on which $S_{4}$ acts, we find that $S_{4}$ is a semidirect product $V\rtimes S_{3},$ as required. (For $n\geq 5,$ Rhodes-primeness of $S_{n}$ with respect to direct products is also an instance of Proposition 6.6(i).)
In case (ii) above, we again have a semidirect product decomposition. To see that $\mathbb{Z}_{p}\rtimes A$ is monolithic with monolith $\mathbb{Z}_{p},$ note that any $g\in\mathbb{Z}_{p}\rtimes A$ that is not in $\mathbb{Z}_{p}$ acts nontrivially on $\mathbb{Z}_{p},$ hence a commutator of $g$ with a nonidentity member of $\mathbb{Z}_{p}$ is a nonidentity element of $\mathbb{Z}_{p},$ so no such $g$ can belong to a normal subgroup not containing $\mathbb{Z}_{p}.$
In case (iii), Rhodes-primeness with respect to direct products is Proposition 6.6(iii). ∎
Each of the above examples shows that the first implication of Lemma 6.3 is not reversible. I was unsure whether the second of those implications might be reversible, but Alexander Olshanskiy provided Example 6.8 below, showing that for every prime $p$ there is a group of order $p^{5}$ for which that reverse implication fails, and Example 6.9, showing that for $p=2,$ there are also two such examples of order $p^{3}.$ Example 6.10 and the paragraph following it give examples of order $p^{3}$ for all odd primes $p.$ The referee observes that all these are what are called “extraspecial p\!p-groups”, namely, p\!p-groups $G$ whose center is isomorphic to $\mathbb{Z}_{p},$ and such that the quotient of $G$ by its center has exponent $p$ [13].
Example 6.8 (A. Olshanskiy).
Let $p$ be any prime, and $H$ the group (of order $p^{3})$ of upper triangular $3\times 3$ matrices over $\mathbb{Z}_{p}$ with diagonal $I.$ Within $H,$ let $a=I+e_{23},$ $b=I+e_{12},$ $c=I+e_{13}.$ Let $G$ (of order $p^{5})$ be the quotient of $H\times H$ by the subgroup generated by the central element $(c,\,c^{-1}).$
Then $G$ is monolithic, with monolith the subgroup $M$ generated by the image of $(c,I),$ which is also the image of $(I,c).$
By construction, $G$ is a quotient, hence a subquotient, of $H\times H;$ but $G$ is not a subquotient of $H,$ since $H$ has smaller order. Hence $G$ is not Rhodes-prime with respect to direct products in the category of finite groups.
Proof.
In $H,$ we find that $a^{p}=b^{p}=c^{p}=I,$ that $c$ is central, and that $ba=abc.$ These properties allow us to write every element in the form $a^{i}\,b^{j}\,c^{k}$ $(0\leq i,j,k<p),$ and since the total number of such expressions is $p^{3},$ the order of $H,$ this expression for each element must be unique. It is easy to verify that the subgroup generated by $c$ is the commutator subgroup of $H,$ the center of $H,$ and a monolith in $H.$
In $G,$ the subgroup $M$ described in the second paragraph of the example is central, hence normal, and has order $p,$ hence is minimal. To see that it is the only minimal normal subgroup, note that any $g\in G$ that is not in $M$ is the image of an element $(h_{0},h_{1})\in H\times H$ such that at least one of $h_{0}$ or $h_{1}$ involves a nonidentity power of $a$ or of $b.$ Assuming without loss of generality that $h_{0}$ involves a nonidentity power of $a,$ we find that the commutator of $(h_{0},h_{1})$ with $(b,I)$ is a nonidentity power of $(c,I).$ Hence in $G,$ the commutator of $g$ with the image of $(b,I)$ is a generator of $M.$ Thus, every normal subgroup of $G$ not contained in $M$ contains $M,$ so as $M$ is simple, $G$ is indeed monolithic with monolith $M.$
The assertions of the final paragraph are immediate. ∎
Example 6.9 (A. Olshanskiy).
Let us write $Q_{8}$ for the quaternion group $\{\pm 1,\,\pm i,\,\pm j,\,\pm k\},$ and $D_{4}$ for the dihedral group (the symmetry group of the square, $p,\,q\,\,|\,\,p^{4}=e,\,\,q^{2}=e,\,\,q\,p\,q^{-1}=p^{-1}\!\!>),$ both of order $8.$
Then $Q_{8}$ is a subquotient of $D_{4}\times D_{4},$ and $D_{4}$ a subquotient of $Q_{8}\times Q_{8},$ but neither $Q_{8}$ nor $D_{4}$ is a subquotient of the other, hence neither group is Rhodes-prime with respect to direct products in the category of finite groups.
However, each is monolithic, with monoliths $\{\pm 1\}\lhd Q_{8}$ and $\{e,\,p^{2}\}\lhd D_{4}$ respectively.
Proof.
We shall show that the 16\!16\!-element group
(6.9) $A\,=~<\!x,\,y\,\,|\,\,x^{4}=e,\,\,y^{4}=e,\,\,y\,x\,y^{-1}=x^{-1}\!>$
(a semidirect product $x\,|\,x^{4}=e\!>\rtimes<\!y\,\,|\,\,y^{4}=e\!>)$ is isomorphic both to a subgroup of $Q_{8}\times Q_{8}$ and to a subgroup of $D_{4}\times D_{4},$ and has both $Q_{8}$ and $D_{4}$ as homomorphic images, from which the above subquotient assertions follow.
Within $Q_{8}\times Q_{8},$ let $x=(i,1)$ and $y=(j,j).$ It is easy to check by looking at first coordinates, and then at second coordinates, that $x$ and $y$ satisfy the relations of (6.9), hence the group they generate is a homomorphic image of $A.$ That group admits a homomorphism onto $Q_{8},$ given by projection to the first component; but $x^{2}\,y^{2}=(1,-1)$ is in the kernel of that homomorphism, hence the group generated by $x$ and $y$ must have larger order than $Q_{8};$ hence it can’t be a proper homomorphic image of the 16\!16\!-element group  $A,$ so it must be isomorphic thereto.
Likewise, within $D_{4}\times D_{4},$ consider the subgroup generated by $x=(p,1)$ and $y=(q,p).$ As in the preceding paragraph, we verify that this group is a homomorphic image of $A,$ and that the projection onto the first component maps it surjectively to $D_{4}.$ However, $y^{2}=(e,p^{2})$ is a nonidentity element of the kernel of that homomorphism, so again the whole group is of larger order than $Q_{8},$ and so must be isomorphic to $A.$
The final monolithicity assertions are easily verified. ∎
(We remark that $D_{4}$ is a semidirect product, $p\,\,|\,\,p^{4}=e\!>\,\rtimes<\!q\,\,|\,\,q^{2}=e\!>\!.$ However, $p\,\,|\,\,p^{4}=e\!>$ is not the least nontrivial normal subgroup $M$ of $D_{4}$ – that is a subgroup thereof – so the above non-Rhodes-primeness result does not contradict Proposition 6.6(ii).)
Here is one more example, which answers a question raised by the referee:
Example 6.10.
For any odd prime $p,$ the monolithic group $G=\mathbb{Z}_{p^{2}}\rtimes\mathbb{Z}_{p}$ (where a generator of $\mathbb{Z}_{p}$ acts on $\mathbb{Z}_{p^{2}}$ by multiplication by 1+p,1\!+\!p, and the monolith is the subgroup $p\,\mathbb{Z}_{p^{2}}\rtimes\{e\})$ is isomorphic to a subquotient of $\mathbb{Z}_{p^{2}}\times H$ where, as in Example 6.8, $H$ is the group of upper triangular $3\times 3$ matrices over $\mathbb{Z}_{p}$ with diagonal $I.$ Hence $G$ is not Rhodes-prime with respect to direct products in the category of finite groups.
Sketch of proof.
In $\mathbb{Z}_{p^{2}}$ we will use additive notation, writing $\mathbb{Z}_{p^{2}}=\{[0],[1],\dots,\,[p^{2}{-}1]\},$ while in $H$ we will use the multiplicative notation of Example 6.8, writing $H=\{a^{i}\,b^{j}\,c^{k}\ |\ 0\leq i,\,j,\,k<p\}$
Within $\mathbb{Z}_{p^{2}}\times H,$ take the subgroup of elements such that the $\mathbb{Z}_{p^{2}}\!-$coordinate, and the exponent of $a$ in the H\!H\!-coordinate, agree modulo $p;$ and divide this subgroup by the congruence equating the central elements ([0],c)([0],\,c) and ([p],I)([p],\,I) (both of order $p).$
The resulting subquotient is a group of order $p^{3}$ in which the image of $([1],\,b)\in\mathbb{Z}_{p^{2}}\times H$ generates a subgroup of order $p^{2},$ and the image of ([0],a)([0],\,a) generates a subgroup of order $p,$ conjugation by which sends the generator of the former subgroup to its 1+p\!1{+}p-th power; so this subquotient is isomorphic to $G=\mathbb{Z}_{p^{2}}\rtimes\mathbb{Z}_{p}.$ But $G$ is clearly not isomorphic to a subquotient of either $\mathbb{Z}_{p^{2}}$ or $H.$
Yet another example is the group denoted $H$ in Examples 6.8 and 6.10 above. That this monolithic group is not Rhodes-prime follows from [3, Theorem 5.2 (ii), p. 204] (where $M$ denotes the group we are calling $H,$ and $N$ the group called $G$ in Example 6.10 above). Namely, the case $k=r=2$ of that statement says that the quotient group of $G\times G$ obtained by identifying the monoliths of the two factors is isomorphic to the quotient group of $G\times H$ obtained in the same way. Hence $H$ is a subquotient of $G\times G;$ but since $G$ and $H$ have the same order, $H$ is not a subquotient of $G,$ proving it non-Rhodes-prime with respect to direct products.
These examples lead one to wonder
Question 6.11.
For $p$ a prime, what finite p\!p\!-groups are Rhodes-prime with respect to direct products in the category of finite groups? Are there any other than the cyclic p\!p\!-groups $\mathbb{Z}_{p^{n}}$?
On the other hand, are there any finite monolithic groups not Rhodes-prime with respect to direct products that aren’t p\!p\!-groups? Yes. Modifying a construction suggested by the referee, let us consider the variant of Example 6.8 gotten by taking $p>2,$ and letting $H$ consist of all upper triangular $3\times 3$ matrices over $\mathbb{Z}_{p}$ in which $e_{11}$ and $e_{33}$ have coefficient $1,$ but the coefficient of $e_{22}$ can be any nonzero member of $\mathbb{Z}_{p};$ thus $|H|=(p{-}1)\,p^{3}.$ Obtain $G,$ as before, by dividing $H\times H$ by the normal subgroup of elements $(I+fe_{1,3},\,I-fe_{1,3})$ for $f\in\mathbb{Z}_{p};$ so $|G|=(p{-}1)^{2}\,p^{5}.$ The reader can verify that the image of the subgroup of elements $(I+fe_{1,3},\,I)$ is again a monolith. As before, though $G$ is a subquotient of $H\times H,$ it is too large to be a subquotient of $H,$ hence it is not Rhodes-prime with respect to direct products. (One can generalize this construction in a way that allows the monolith to have order $2$: Replace $\mathbb{Z}_{p}$ by any finite field of cardinality >2,>2, in particular allowing finite proper extensions of $\mathbb{Z}_{2},$ and at the end of the construction, divide the central subgroup of elements $(I+fe_{1,3},\,I+g\,e_{1,3})$ not just by the subgroup of elements $(I+fe_{1,3},\,I-fe_{1,3}),$ but by a large enough overgroup thereof to shrink it down to a cyclic group, so that it again becomes a monolith.)
Moving on, it is easy to give examples showing nonreversibility of the last implication of Lemma 6.3. E.g.:
Example 6.12.
Let $p$ be any prime, and $q_{0},$ $q_{1}$ primes (possibly equal to one another) that are both $\equiv 1\pmod{p}$, so that the automorphism groups of $\mathbb{Z}_{q_{0}}$ and $\mathbb{Z}_{q_{1}}$ both have orders divisible by $p.$ Thus, $\mathbb{Z}_{p}$ has faithful actions on both these groups. Let $G=(\mathbb{Z}_{q_{0}}\times\mathbb{Z}_{q_{1}})\rtimes\mathbb{Z}_{p},$ defined using these actions.
Then $G$ is Tarski-prime in the category of finite groups, but is not monolithic.
Proof.
To show Tarski-primeness, note that the order of $G,$ $p\,\,q_{0}\,q_{1},$ is a product of just three primes, hence if $G$ were a nontrivial direct product, one of the factors would have prime order, hence be commutative, and since it centralizes the other factor, it would be central in $G.$ But $G$ has no nonidentity central elements; so it is not such a product, so by the Krull-Schmidt Theorem, $G$ is Tarski-prime in the category of finite groups.
But $G$ is not monolithic, since it has at least the two minimal normal subgroups $\mathbb{Z}_{q_{0}}$ and $\mathbb{Z}_{q_{1}}.$ (These are its only minimal normal subgroups if $q_{0}\neq q_{1}.$ If $q_{0}=q_{1},$ on the other hand, then $\mathbb{Z}_{q_{0}}\times\mathbb{Z}_{q_{1}}$ is a 2\!2\!-dimensional $\mathbb{Z}_{q_{0}}\!$-vector space, and has $q_{0}\!+\!1$ one-dimensional subspaces; and if we give $\mathbb{Z}_{p}$ the same action on both of those copies of $\mathbb{Z}_{q_{0}},$ then all $q_{0}\!+\!1$ of those subspaces are minimal normal subgroups of $G.)$
The examples and proofs given above make much use of subgroups, but less of homomorphic images. This led me to wonder: Suppose we call an object modified Rhodes-prime with respect to direct or semidirect products if it satisfies condition (iv) or (v) of Convention 6.1 with “subquotient” replaced by “subgroup”. How do these conditions compare with the unmodified conditions?
These questions are easily answered for finite groups:
A finite group $G$ is modified Rhodes-prime with respect to direct products if and only if it is monolithic. Indeed, if $G$ has a monolith $M,$ then for any embedding $G\to G_{1}\times G_{2},$ the kernels of the induced maps $\to G_{1}, GG$$\to G_{2}$, having trivial intersection, cannot both contain $M,$ so one of them must be trivial, giving an embedding of $G$ in $G_{1}$ or $G_{2},$ proving modified Rhodes-primeness. On the other hand, any non-monolithic finite group $G$ has nontrivial normal subgroups $N_{1},\,N_{2}$ with trivial intersection, giving an embedding $G\to G/N_{1}\times G/N_{2},$ and looking at cardinalities, we see that $G$ cannot be embedded in either $G/N_{1}$ or $G/N_{2},$ so $G$ is not modified Rhodes-prime with respect to direct products.
On the other hand, we can see from the proofs of Propositions 6.4 and 6.5 that modified Rhodes-primeness of finite groups with respect to semidirect products is equivalent to simplicity, hence equivalent to ordinary Rhodes-primeness with respect to semidirect products.
Note that the modified Rhodes-primeness conditions are not, a priori, either stronger or weaker than the unmodified conditions: The implication defining each of the modified conditions has a strengthened hypothesis, that $G$ be isomorphic to a subgroup of a group $G_{0}\times G_{1}$ or $G_{0}\rtimes G_{1},$ not merely to a subquotient; but it has a similarly strengthened conclusion. For finite groups, the modified Rhodes-primeness conditions turned out to be weaker than or equivalent to the unmodified conditions, but for not-necessarily-finite groups, or various classes thereof, the relationships may be different.
If we look at the condition of modified Rhodes-primeness with respect to direct products on not necessarily finite groups, a notable class of examples are the infinite abelian groups of prime exponent $p.$ If such a group $G$ is embedded in a direct product $G_{1}\times G_{2},$ then regarding $G$ as a $\mathbb{Z}_{p}\!-$vector space, its images in $G_{1}$ and $G_{2}$ will also be $\mathbb{Z}_{p}\!-$vector spaces, at least one of which must have the same infinite dimension as $G,$ hence $G$ can be embedded in one of those groups (even if not via the given map to that group). Thus $G$ is modified-Rhodes prime with respect to direct products. But $G$ does not have a minimal normal subgroup, or even satisfy the condition that the intersection of any two nontrivial normal subgroups be nontrivial – it has many pairs of nonzero $\mathbb{Z}_{p}\!-$subspaces with zero intersection. I have not examined whether $G$ also has the unmodified Rhodes-primeness condition with respect to direct products, and/or the modified or unmodified condition with respect to semidirect products.
Returning to the unmodified conditions, another question I have not thought hard about is
Question 6.13.
If a finite group $G$ is Rhodes-prime with respect to direct products in the category of finite groups, must it also have that property in the category of all groups? (The converse is clearly true.)
Propositions 6.4 and 6.5 show that the answer to the corresponding question for Rhodes-primeness with respect to semidirect products is “yes”, and (4.5) shows the same for Tarski-primeness.
Some further observations:
The analog of the Krull-Schmidt theorem is not true for finite semigroups. To see this, first note that in the proof of Lemma 2.5 above, the case of finite $\kappa$ uses only finite semigroups, and so shows that $\textup{Null}(\kappa)$ is not Tarski-prime among finite semigroups. Taking $\kappa$ such that $\kappa\!+\!1$ is a prime number, $\textup{Null}(\kappa)$ also cannot be a nontrivial direct product. For another example see [9, Exercise 4, p. 265].
But as we saw in §§ 3-4, things become much better if we move from semigroups to monoids. In particular, the Jónsson-Tarski Theorem ([7], [9, p. 290]) proves unique factorization for finite algebras in a large class of varieties including the variety of monoids, and hence Tarski-primeness of all nontrivial non-factoring algebras in those varieties. (Regarding the language used in [9], see discussion preceding and following (4.5) above. That discussion concerns a generalization, [9, Exercise 5.15 (10), p. 300], of the Jónsson-Tarski Theorem to not necessarily finite algebras whose “centers”, in the sense used there, are finite.)
These thoughts lead to our last question.
Question 6.14.
Understanding the two sorts of Rhodes-primeness to be defined for semigroups and monoids as we define them in Convention 6.1 for groups (indeed, they are so defined for semigroups in [11]), and likewise for the condition of Tarski-primeness (which is defined for general algebras in [9]), the following five implications are clear. Are some or all of them reversible?
For any finite group $G,$
(6.10) $G$ is Rhodes-prime with respect to semidirect products as a finite semigroup $\implies$ $G$ is Rhodes-prime with respect to semidirect products as a finite monoid $\implies$ $G$ is Rhodes-prime with respect to semidirect products as a finite group.
(6.11) $G$ is Rhodes-prime with respect to direct products as a finite semigroup $\implies$ $G$ is Rhodes-prime with respect to direct products as a finite monoid $\implies$ $G$ is Rhodes-prime with respect to direct products as a finite group.
(6.12) $G$ is Tarski-prime as a finite semigroup $\implies$ $G$ is Tarski-prime as a finite monoid, equivalently, as a finite group.
Likewise, for arbitrary groups $G,$ are some or all of the corresponding implications with “finite” everywhere deleted reversible?
One can see from the Jónsson-Tarski Theorem mentioned above that a finite group is Tarski-prime as a monoid if and only if it is Tarski-prime as a group, which is why (6.12) above shows fewer conditions than the other two displays. (But in the final sentence of Question 6.14 the analog of (6.12) could be expanded to a 3\!3\!-condition implication like the others.)
Incidentally, because a semigroup or monoid is not in general isomorphic to its opposite, there are actually two versions of the concept of semidirect product for these objects, one based on “action on the right” and the other on “action on the left”. But since the opposite of every semigroup or monoid is still a semigroup or monoid, general results about each of these constructions imply the corresponding results about the other. Cf. [11, p. 24, first sentence of next-to-last paragraph].
7. An observation on Rhodes-primeness with respect to direct
products in arbitrary varieties of algebras
The concept of Rhodes-primeness with respect to direct products, as defined in Convention 6.1(iv) above, makes sense in any variety of algebras. (In contrast, in a general variety there’s no obvious version of the concept of semidirect product of objects, hence no obvious version of the concept of Rhodes-primeness with respect to semidirect products.) We end this note with a result on finite algebras Rhodes-prime with respect to direct products in general varieties of algebras.
Given an object or family of objects $X$ in a variety $V,$ I will write $\textup{Var}(X)$ for the subvariety of $V$ generated by $X.$ We allow $V$ to have infinitely many operations, and/or to have operations of infinite arities – the finiteness of the algebras considered will make those features unimportant.
Abandoning the convention made for semigroups at the beginning of this note, we will understand every variety with no zeroary operations to have an empty algebra. In such a variety, this will be a subquotient of every object, hence will be considered Rhodes-prime with respect to direct products. We likewise put aside the restriction made in Convention 6.1 of excluding 1\!1\!-element algebras from the classes called Rhodes-prime.
Recall that an element $x$ of a lattice is called join prime if $x\leq y\vee z$ implies $x\leq y$ or $x\leq z.$
Proposition 7.1.
Let $V$ be any variety of algebras (in the sense of universal algebra) and $X$ a finite object of $V.$ Then the following conditions are equivalent:
(i)$X$ is Rhodes-prime with respect to direct products in the category of finite objects of $V.$
(ii)  If $Y_{0}$ and $Y_{1}$ are finite nonempty objects of $V$ such that $X\in\textup{Var}(\{Y_{0},\,Y_{1}\}),$ then $X$ is a subquotient either of $Y_{0}$ or of $Y_{1}.$
(iii) $\textup{Var}(X)$ is join-prime in the join-semilattice of subvarieties of $V$ generated by finite families of finite algebras; and whenever $Y$ is a finite nonempty algebra in $V$ such that $\textup{Var}(X)\subseteq\textup{Var}(Y),$ $X$ is a subquotient of $Y.$
Proof.
[…]
Assuming (i), suppose we are given $Y_{0},$ $Y_{1}$ as in the hypothesis of (ii). Then taking $n$ such that $X$ is generated by $n$ elements, the assumption $X\in\textup{Var}(\{Y_{0},\,Y_{1}\})$ makes $X$ a homomorphic image of the free algebra on $n$ generators in $\textup{Var}(\{Y_{0},\,Y_{1}\}),$ which can be constructed as a subalgebra of $Y_{0}^{\textup{card}(Y_{0})^{n}}\!\!\!\times Y_{1}^{\textup{card}(Y_{1})^{n}}\!\!.$ This makes $X$ a subquotient of a finite direct product of copies of $Y_{0}$ and $Y_{1}.$ Applying the Rhodes-primeness condition (i) inductively, we conclude that that $X$ is a subquotient of $Y_{0}$ or of $Y_{1},$ proving (ii).
Now assume (ii). Since any subvariety of $V$ generated by finitely many finite algebras is generated by one finite algebra (the direct product of the nonempty algebras in the family), the first assertion of (iii) is equivalent to saying that if $\textup{Var}(X)\subseteq\textup{Var}(\{Y_{0},\,Y_{1}\}),$ then $\textup{Var}(X)\subseteq\textup{Var}(Y_{0})$ or $\textup{Var}(X)\subseteq\textup{Var}(Y_{1}).$ This condition is trivial if one of $Y_{0},$ $Y_{1}$ is empty, while assuming them nonempty, it implied by the subquotient conclusion of (ii). The second assertion of (iii) is obtained from (ii) by taking $Y_{0}=Y_{1}=Y.$
Finally, assuming (iii), suppose $X$ is a subquotient of $Y_{0}\times Y_{1}.$ Then in particular, it belongs to the variety generated by $Y_{0}$ and $Y_{1},$ so by the first condition of (iii), $X$ belongs to the variety generated by one of these, which we will call $Y.$ Then assuming $Y$ nonempty, the second condition of (iii) shows that $X$ is a subquotient of $Y,$ establishing (i). On the other hand, if $Y$ is empty, then so is $X,$ and since the empty algebra (if it exists in $V)$ is Rhodes-prime with respect to direct products, we again get (i). ∎
We note that neither of the two conditions of (iii) implies the other. Indeed, let $V$ be the variety of abelian groups. Note that every subvariety of $V$ is determined by an identity $x^{n}=e$ for a unique nonnegative integer $n.$ It is easy to see that for any prime $p,$ the group $X=\mathbb{Z}_{p}\times\mathbb{Z}_{p}$ satisfies the first part of (iii) but (taking $Y=\mathbb{Z}_{p}),$ not the second, while for any two distinct primes $p\neq q,$ the group $X=\mathbb{Z}_{p}\times\mathbb{Z}_{q}$ satisfies the second but not the first.
In that same condition (iii), can the inclusion $\textup{Var}(X)\subseteq\textup{Var}(Y)$ be replaced by $\textup{Var}(X)=\textup{Var}(Y)$? Here is a version of an example provided by the referee showing that it cannot. Let $G$ be a nontrivial finite group, $V$ the variety of G\!G\!-sets, and $X$ a 2\!2\!-element set on which $G$ acts trivially. Thus, $X$ is a subquotient of a G\!G\!-set $Y$ if and only if $Y$ has more than one orbit. In particular, if $Y$ is a free G\!G\!-set on one generator, then $X$ is not a subquotient of $Y,$ but is a subquotient of $Y\times Y;$ so $X$ is not Rhodes-prime with respect to direct products. But turning to condition (iii), we see that $\textup{Var}(X)$ is the variety of G\!G\!-sets on which $G$ acts trivially, which is a minimal nontrivial element of the lattice of subvarieties of $V,$ hence join-prime; and we easily see that the “whenever” condition of (iii) with $\subseteq$ replaced by == also holds for $X;$ so that modified version of (iii) is not equivalent to (i).
8. Acknowledgements
I am grateful to Jonathan Farley for his group-email pointing out that the result of [14] answers one of the questions listed at [9, p. 263], which drew my attention to that list of questions, in particular the one now answered in §2 above; to Ralph Freese and Walter Taylor for information about what is known about Tarski-primeness; to Benjamin Steinberg, Stuart Margolis and Edmond Lee for information about what is known about the two Rhodes-primeness conditions; to Alexander Olshanskiy for Proposition 6.5, Examples 6.8 and 6.9, and some pointers on terminology, and to the referee for further information about known results and terminology, and some useful examples and suggestions.